OpenStudy (anonymous):

Always reward medals, dy/dx=ysin(2x) when x=0 y=3

OpenStudy (gorv):

dy/dx=y*sin2x \[\int\limits_{}^{} dy/y=\int\limits_{}^{}\sin2x*dx\]

OpenStudy (gorv):

lny=(-cos2x)/2+c now at x=0 y=3 ln3=(-cos0)/2+c c=ln3 +1/2 lny=-cox2x/2+ln3+1/2 lny-ln3=-cos2x/2+1/2=(1-cos2x)/2 ln(y/3)=(1-cos2x)/2 y/3=e^((1-cos2x)/2) y=(1/3)*e^((1-cos2x)/2)

OpenStudy (gorv):

@hhuumm

OpenStudy (anonymous):

answer key states y=3e^[sin^2(x)]

OpenStudy (anonymous):

how did you get cos2x/2? integral of sin2x= -cos2x/2?

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

yep

OpenStudy (anonymous):

ok stick around just a little bit longer so i can wrap my head around it...

OpenStudy (gorv):

pls hurry

OpenStudy (anonymous):

lol trust me my exam is in 40 minutes i'm moving

OpenStudy (anonymous):

instead of adding ln and blah blah could i have just put both sides as an exponent to the E and called it a day (step 4)

OpenStudy (gorv):

try to put it u will get ur ans

OpenStudy (anonymous):

despite the awk glamour shot avi, you are the man. Earned your medal thanks.